In choosing between four new jobs, a business graduate considers the potential payoffs over the next three years. The following table contains four alternative organizations and two states of nature represented by slow promotion and fast promotion. The payoffs for each alternative and each sate of nature is given is the table below.

Alternative

Slow Promotion

Fast Promotion

H. High-flying consultant

($180,000)

$600,000

U. Utility analyst

$200,000

$400,000

R. Research assistant

$250,000

$260,000

I. IT analyst

$240,000

$320,000

Which alternative would be chosen using maximin decision rule?

a.         H.

b.         U.

c.         R.

d.         I.

Which alternative would be chosen using maximax decision rule?

a.         H.

b.         U.

c.         R.

d.         I.

Which alternative would be chosen using Laplace decision rule?

a.         H.

b.         U.

c.         R.

d.         I.

Which alternative would be chosen using minimax regret decision rule?

a.         H.

b.         U.

c.         R.

d.         I.
Answer #1

(1)

Maximin decision rule:

In maximin, the individual chooses the best option from the lowest individual payoff. For each alternative, find the minimum pay off, then select the maximum payoff from the calculated minimum payoff.

Alternative

Slow Promotion

Fast Promotion

H

($180,000)

$600,000

($180,000)

U

$200,000

$400,000

$200,000

R

$250,000

$260,000

$250,000

I

$240,000

$320,000

$240,000

The correct answer: Option (c) – R

===============================================================

(2)

Maximax decision rule:

In this decision rule, the best among the best alternative is chosen. First, find the best payoff from each alternatives, then select the maximum payoff from the identified payoffs.

Alternative

Slow Promotion

Fast Promotion

H

($180,000)

$600,000

$600,000

U

$200,000

$400,000

$400,000

R

$250,000

$260,000

$260,000

I

$240,000

$320,000

$320,000

The correct answer: option (a) – H

===============================================================

(3)

Laplace decision rule:

According to Laplace rule, the equal probability of occurring each state of nature is assigned and then the overall payoff is calculated.

Alternative

Slow Promotion

Fast Promotion

H

($180,000)

$600,000

(-180,000*0.5) +(600,000*0.5) = $210,000

U

$200,000

$400,000

(200,000*0.5) +(400,000*0.5) = $300,000

R

$250,000

$260,000

(250,000*0.5) +(260,000*0.5) = $255,000

I

$240,000

$320,000

(240,000*0.5) +(320,000*0.5) = $280,000

The correct answer: option (b) – U

===============================================================

(4)

Minimax regret decision rule calculates the regret for each payoff by subtracting the individual payoff from the best payoff. From the calculated regret the minimum regret is selected.

Regret table and payoff table:

Alternative

Slow Promotion

Fast Promotion

H

600,000 – (-180,000) = $780,000

600,000 – 600,000 = $0

$780,000

U

400,000-200,000 = $200,000

400,000 – 400,000 = $0

$200,000

R

260,000 – 250,000 = $10,000

260,000 – 260,000 = $0

$10,000

I

320,000 – 240,000 = $80,000

320,000 – 320,000 = $0

$80,000

The correct answer: option (c) - R

Please upvote, if you find the answer helpful. It helps us to improve quality and other students to identify the best answer.